Zachary Company sells lamps and other lighting fixtures. The purchasing department manager prepared the following inventory purchases budget. Zachary’s policy is to maintain an ending inventory balance equal to 20 percent of the following month’s cost of goods sold. April’s budgeted cost of goods sold is $83,000.

Required

  1. Complete the inventory purchases budget by filling in the missing amounts.

  2. Determine the amount of cost of goods sold the company will report on its first quarter pro forma income statement.

  3. Determine the amount of ending inventory the company will report on its pro forma balance sheet at the end of the first quarter.

  4. Inventory Purchases Budget
    January February March
    Budgeted cost of goods sold $54,000 $58,000 $64,000
    Plus: Desired ending inventory 11,600
    Inventory needed 65,600
    Less: Beginning inventory 10,800
    Required purchases (on account) $54,800

    Complete the inventory purchases budget by filling in the missing amounts.

    Determine the amount of cost of goods sold the company will report on its first quarter pro forma income statement.Determine the amount of ending inventory the company will report on its pro forma balance sheet at the end of the first quarter.

    b. Cost of goods sold
    c. Ending inventory
Answer #1
A
Inventory Purchases Budget
January February March
Budgeted cost of goods sold 54000 58000 64000
Plus: Desired ending inventory 11600 12800 16600
Inventory needed 65600 70800 80600
Less: Beginning inventory 10800 11600 12800
Required purchases (on account) 54800 59200 67800
B
Cost of goods sold 176000 =54000+58000+64000
C
Ending inventory 16600 =83000*20%
Workings:
January February March
Plus: Desired ending inventory =58000*20% =64000*20% =83000*20%